# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v1

@@ I Introduction
In the realm of computational finance, it has become nearly a conventional practice to convert raw historical asset information into market trend indicative signals, referred to as alpha factor values[1]. The functions that generate these signals are known as alpha factors, which can be expressed in two distinct forms: the deep model and the formula[2, 3, 4]. Mining high-quality alpha factors has become a trendy topic among investors and researchers, due to their relationship with excess investment returns.
@@
Comparatively, alpha factors represented in formulaic forms enjoy much better interpretability and thus are favored by market participants. The most intuitive method for automatically mining formulaic factors is the tree-based models, which mutate expression trees to generate new alpha factors[6, 7, 8]. Another commonly used method is the genetic programming, which generates the formulaic alpha factor expressions through iterative simulations of genetic variation and natural selection in biological evolution[9, 10, 11].
@@
Unfortunately, tree-based models and genetic programming still face certain challenges. Tree-based models, while easy to understand and implement, may encounter performance bottlenecks when dealing with non-linear relationships and high-dimensional data[12]. Genetic programming, on the other hand, can deal with broader types of expressions by properly defining the search space, including complex nonlinear and non-parametric factor expressions, but it usually fails to explore the search space of large-scale expressions. Additionally, the computational costs of genetic programming are usually expensive[13].
@@
Recently, Yu et. al. [14] proposes a new framework for mining formulaic alpha factors with deep reinforcement learning, trying to bridge both deep learning and formulaic based methods. It employs Markov Decision Processes (MDPs)[15] to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using Deep Reinforcement Learning (DRL). It essentially aims at finding formulaic alpha factors. With the help of DRL, it also overcomes the limitations of the explorative search ability suffered by traditional tree models and genetic programming[14].
The key to this framework lies in using the Reverse Polish Notation (RPN) to convert formulaic alpha factors into a sequence composed of tokens. Tokens include various operators, the original asset features, constants, and other elements that can constitute a formula, which is the action in the MDPs, with the state being a sequence composed of the previous tokens. In this framework, once a new action is selected, the new state is uniquely determined, as the next step sequence is composed of the present step sequence and the newly generated token. Moreover, only after a complete sequence has been generated can the performance of this factor be evaluated, i.e., rewards are obtained, giving the MDPs of this framework the characteristic of sparse rewards.
[14] employed the Proximal Policy Optimization (PPO) algorithm, which is a natural solution based on the actor-critic framework, using a value network to reduce the variance in the training process. Although [14] has addressed the issues with traditional methods, we believe that PPO is not suitable for this factor mining framework. The aforementioned framework has the characteristics of deterministic state transitions (transition function satisfies the Dirac distribution) and trajectory-level-reward (non-zero reward can only be obtained after completing a full trajectory), but PPO, due to the inefficiency of the value network in trajectory-level-reward environment[16], and the fact that the value network generally has the same scale as the policy network, leading to twice the time for one iteration of the parameters, cannot effectively solve MDPs for mining formulaic alpha factors.
There are two ways to estimate the policy gradient: Temporal Difference Sampling (e.g., PPO using the actor-critic framework) and Monte Carlo sampling (e.g., REINFORCE without using the actor-critic framework). Due to the lack of the value network, we would like to discard it, and naturally thought of choosing REINFORCE. The REINFORCE method does not rely on immediate rewards but on cumulative rewards, which is better suited for our task. Additionally, when using REINFORCE to solve the problem of formulaic alpha factor mining, the sampling process can be completed very quickly and at a low cost, meaning that the cost of obtaining cumulative returns is not high. This is because, unlike alpha factors based on deep models, once the policy network converges, it only needs to perform a single forward pass to obtain the formulaic alpha factors and quickly calculate the factor values for each time’s assets, whereas the former requires a forward pass for each factor value calculation, which brings a significant computational resource consumption. However, REINFORCE also suffers from high variance, resulting in a training curve that often fails to converge.
Based on the discussions mentioned above, our work proposes a new reinforcement learning method outside the actor-critic framework for the first time to accomplish the task of formulaic alpha factors mining. This also offer a new perspective on alpha factors mining within the context of reinforcement learning. Compared to the PPO algorithm used in work[14], the time required for our algorithm to converge is significantly reduced and the factors generated are steadier. Our QuantFactor REINFORCE (QFR) algorithm implements a constrained sequence generator to ensure effective formulaic alpha factors generation. By discarding the value network of [14], our algorithm exhibits faster convergence speeds. Since the transition functions in MDPs for formulaic alpha factors mining satisfying Dirac distribution, the high variance issue caused by discarding the value network is mitigated. To further reduce the variance of the policy gradient estimate, QFR employs a greedy policy to generate a novel baseline. This improvements are also theoretically studied. Additionally, we introduce information ration (IR) for reward shaping to better balance returns and risks. With the introduction of a novel baseline, QFR has a reduced and constrained variance compared to REINFORCE. Furthermore, after incorporating the IR testing mechanism, our work not only evaluates the predictive accuracy of factors but also considers the volatility of their predictive signals. It provides a more comprehensive assessment of factor performance and generate steadier factors that can better adapt to changes in market volatility.
@@
To evaluate our proposed QFR algorithm, we conducted extensive experiments on multiple real-world asset datasets. Our experimental results show that the set of formulaic alpha factors generated by the QFR algorithm outperforms those generated by previous methods. It is also shown that theoretical results properly predict trends in the experimental results.
@@
  • We propose a novel, stable and efficient reinforcement learning algorithm for mining formulaic alpha factors. Unlike prior research that relies on the actor-critic framework, we discard the value network. This decision stems from the value network’s inability to effectively manage MDPs with trajectory-level rewards, its equivalent scale to the policy network, which doubles the parameter iteration time. Moreover, since the transition functions in these MDPs satisfying Dirac distribution, the high variance due to the absence of the value network is significantly mitigated.
  • Two new components are proposed in the framework of formulaic alpha factors mining using reinforcement learning. A novel baseline that employs a greedy policy to generate is proposed to reduce the variance of the policy gradient estimate. Additionally, IR is introduced as a reward shaping mechanism to balance returns and risks, providing a more comprehensive assessment of factor performance and generating steadier factors. We also provide a set of theoretical results, which includes an analysis of the training variance under state transition functions with various distributions, the derivation of the upper bound of the variance after introducing the baseline, and the proof that the variance decreases relative to the REINFORCE algorithm.
@@ II-A Automatic Mining for Alpha Factors
The work in [14] is the first to employ reinforcement learning for formulaic alpha factors mining. It employs Markov Decision Processes (MDPs)[15] to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using reinforcement learning. Compared to works [26, 27, 28] that directly use reinforcement learning to build trading agents, this framework uses historical quantitative and price data of assets as input to find a set of formulaic alpha factors with strong interpretability, avoiding the black-box problem and overcoming the limitations of [6, 7, 8, 9, 10] in independently mining individual formulaic factors, such as homogenization among factors, lack of synergy, and difficulty in adapting to dynamic market changes. The characteristics of reinforcement learning, supervised learning, tree models, and heuristic algorithms when applied to mining alpha factors are shown in Table I.
II-B REINFORCE algorithm
III Problem Formulation and Preliminaries
This section introduces the considered formulaic alpha factors and their mining processes. We have modeled a finite-length MDPs for these processes, where a token is generated from a limited token pool at each step. Tokens include various operators, the original volume-price features, fundamental features, time deltas, and constants. The two classical policy gradient algorithms: REINFORCE and PPO can be used as solutions to these MDPs where only the last step reward is non-zero. This kind of reinforcement learning algorithm requires a large number of interactions with these MDPs for formulaic alpha factor mining to optimize the prediction of the next token. Overall, we consider MDPs with only trajectory-level rewards and introduce their nature solutions.
III-A Formulaic Alpha Factors for Predicting Asset Prices
Consider a real market with $n$ assets over $T$ trading days. On each trading day $t\in\{1,2,\cdots,T\}$, each asset $i$ is represented by a feature vector $\mathbf{x}_{ti}\in\mathbb{R}^{m\times\tau}$. This vector consists of $m$ raw market features, such as open, high, low, close, and volume values, over the recent $\tau$ days. Here, $\mathbf{x}_{tij}\in\mathbb{R}^{\tau}$ denotes the sequence of values for the $j$-th raw feature over these $\tau$ days.
Next, we define an alpha factor function $f$, which transforms the feature matrix for all assets on a given trading day into alpha values, represented as $\mathbf{X}_{t}=[\mathbf{x}_{t1},\mathbf{x}_{t2},\cdots,\mathbf{x}_{tn}]^{T}\in\mathbb{R}^{n\times m\times\tau}$, into alpha factor values. Specifically, $f$ maps $\mathbf{X}_{t}$ to a vector $\mathbf{z}_{t}=f(\mathbf{X}_{t})\in\mathbb{R}^{n}$, where $\mathbf{z}_{t}$ holds the alpha factor values for each asset on trading day $t$. The real asset feature dataset $\mathcal{X}=\{\mathbf{X}_{t}\}$.
To measure the effectiveness of a formulaic alpha factor, we need a real asset price dataset $\mathcal{Y}=\{\mathbf{y}_{t}\}$, where $t\in\{1,2,\cdots,T\}$ and $\mathbf{y}_{t}$ is the 5-day asset return, where $\mathbf{y}_{t}\in\mathbb{R}^{n}$. The newly generated formulaic alpha factor will be added to an alpha factors pool, and the combined factor value $\mathbf{z}_{t}^{\prime}$ will be calculated through a combination model, which will be discussed in detail in Section IV-A. The expression for the Pearson correlation coefficient between $\mathbf{y}_{t}$ and the combination factor value $\mathbf{z}_{t}^{\prime}$, also known as the Information Coefficient (IC), is as follows:
$IC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)=\frac{\text{Cov}(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t})}{\sigma_{\mathbf{z}^{\prime}_{t}}\sigma_{\mathbf{y}_{t}}}.$ (1)
The average IC values over all trading days are denoted as $\overline{IC}=\mathbb{E}_{t}\left[IC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)\right]$.
Formulaic alpha factors are mathematical expressions that can be represented using RPN, which is a sequence of tokens. Tokens include various operators, the original volume-price features, fundamental features, time deltas, and constants. The operators include elementary functions that operate on single-day data, known as cross-sectional operators (e.g., Abs($x$) for the absolute value $|x|$, Log($x$) for the natural logarithm log($x$)), as well as functions that operate on a series of daily data, known as time-series operators (e.g., Ref($x$,$t$) for the expression $x$ evaluated at $t$ days before the current day).
Such formulas can naturally be represented by an expression tree, with each non-leaf node representing an operator, and the children of a node representing the original volume-price features, fundamental features, time deltas, and constants being operated on. Each such expression has a unique post-order traversal, using RPN. Examples of some formulaic alpha expressions derived from Alpha101 [29] and their corresponding RPN representations are shown in Table II. An example of a formulaic alpha expression, together with its corresponding tree and RPN representation, is shown in Fig. 1.
@@ III-A Formulaic Alpha Factors for Predicting Asset Prices
Reinforcement Learning Supervised Learning Tree Models Heuristic Algorithms
@@
Figure 1: Three forms of an alpha factor: (a) formulaic alpha expression; (b) tree structure; (c) RPN representation. The Begin (BEG) token and Separator (SEP) token of the RPN representation are used to mark the beginning and end of the Markov Decision Process, which is discussed in detail in Section III-B.
III-B MDPs for Mining Formulaic Alpha Factors
@@
We can model the process of generating a linear sequence that can equivalently represent a formulaic alpha factor as an MDP, which can be described as $\{\mathcal{S},\mathcal{A},P,r\}$, where $\mathcal{S}$ and $\mathcal{A}$ denote the finite state and action spaces, respectively. $P$ is the transition function defining state transitions, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function assigning values to state-action pairs. Our goal is to obtain the policy $\pi_{\mathbf{\theta}}$, where $\mathbf{\theta}$ is the training parameter. It generates tokens from a finite set of alternative tokens. Each state in the MDPs corresponds to a sequence of tokens denoting the currently generated part of the expression, denoted as $\mathbf{s}_{t}=\mathbf{a}_{1:t-1}=[a_{1},a_{2},\cdots,a_{t-1}]^{T}$, the sampling process is modeled as $a_{t}\sim\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{a}_{1:t-1}\right)$. The action $a_{t}$ is the next token following the currently generated part of the expression.
This process continues until the SEP token is encountered or the maximum length is reached. The initial state is always the BEG token, so a valid state always starts with BEG and is followed by previously chosen tokens. Obviously, any generated sequence cannot be guaranteed to be a legal RPN sequence, therefore we only allow specific actions to be taken in certain states to ensure the correct format of the RPN sequence. When $\mathbf{a}_{1:t-1}$ and $\mathbf{a_{t}}$ are already known, then $\mathbf{s}_{t+1}=\mathbf{a}_{1:t}$ is uniquely determined, which means that the state transition function $P$ satisfies the Dirac distribution:
$P\left(\mathbf{s}_{t+1}\mid\mathbf{a}_{1:t}\right)=\left\{\begin{array}[]{ll}1&\text{ if }\mathbf{s}_{t+1}=\mathbf{a}_{t}\\ 0&\text{ otherwise. }\end{array}\right.$ (2)
@@
The optimization objective in these MDPs framework is to learn a policy $\pi_{\mathbf{\theta}}$ that maximizes the expected cumulative reward over time:
$J(\theta)=\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\mathbf{\theta}}}\left[r\left(\mathbf{a}_{1:T}\right)\right]$ (3)
@@
Non-zero rewards are only received at the final step, which evaluates the quality of a complete formulaic factor expression, not individual tokens:
$r\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)=\left\{\begin{array}[]{ll}0&\text{if }t\neq T\\ r\left(\mathbf{a}_{1:T}\right)&\text{ otherwise, }\end{array}\right.$ (4)
@@
where $r\left(\mathbf{a}_{1:T}\right)=\overline{IC}$. Expressions that are syntactically correct might still fail to evaluate due to the restrictions imposed by certain operators. For example, the logarithm operator token is not applicable to negative values. Such invalidity can not be directly detected. Therefore, these expressions are assigned a reward of $-1$ (the minimum value of IC) to discourage the policy from generating these expressions.
III-C Comparing REINFORCE with PPO
The natural solutions to these MDPs for formulaic alpha factors mining are the respected Proximal Policy Optimization (PPO) algorithm[44], which uses the classic Policy Gradient:

### v2

@@ I Introduction
Constructing a superior portfolio involves distinguishing noise from the huge raw data in financial markets and discovering signals to balance risk and return, which is widely regarded as a signal processing problem[1, 2, 3]. In the realm of computational finance, it has become nearly a conventional practice to convert raw historical asset information into into alpha factor values[4], which serve as indicative signals of market trends for portfolio management[5]. The functions that generate these signals are known as alpha factors, which can be expressed in two distinct forms: the deep model and the formula[6, 7, 8]. Mining high-quality alpha factors has become a trendy topic among investors and researchers, due to their relationship with excess investment returns[9].
@@
Comparatively, alpha factors represented in formulaic forms enjoy much better interpretability and thus are favored by market participants. The most intuitive method for automatically mining formulaic factors is the tree-based models, which mutate expression trees to generate new alpha factors[11, 12, 13]. Another commonly used method is the genetic programming (GP), which generates the formulaic alpha factor expressions through iterative simulations of genetic variation and natural selection in biological evolution[14, 15, 16]. Unfortunately, both methods still face certain challenges. Tree-based models, while easy to understand and implement, may encounter performance bottlenecks when dealing with non-linear relationships and high-dimensional data[17]. Genetic programming, on the other hand, can deal with broader types of expressions by properly defining the search space, including complex nonlinear and non-parametric factor expressions, but it usually fails to explore the search space of large-scale expressions. Additionally, it is usually computationally expensive[18].
@@
Recently, Yu et al. [19] proposes a promising framework named AlphaGen for mining formulaic alpha factors with Deep Reinforcement Learning (DRL), trying to bridge both deep learning and formulaic based methods. It employs Markov Decision Processes (MDPs)[20] to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using DRL. It essentially aims at finding formulaic alpha factors. With the help of DRL, it also overcomes the limitations of the explorative search ability suffered by traditional tree models and genetic programming[19].
The key to this framework lies in using the Reverse Polish Notation (RPN) to convert formulaic alpha factors into a sequence of tokens. Tokens include various calculation operators, the original asset features, financial constants, and other elements that can constitute a formula. The action in the MDPs is represented as a new token to be added to the sequence, and the state of the MDPs is the current sequence composed of the previous tokens (actions). The reward can be set as any well-established performance indicators of alpha factors. The remaining problem is how to learn the policy effectively. There are generally two ways to estimate the policy gradient: Temporal Difference Sampling (e.g., the actor-critic architecture) and Monte Carlo sampling (e.g., without using the critic network). The Temporal Difference method may be biased because it adopts the estimated values of the critic network rather than actual rewards, but its variance can be made small as it updates the networks after each action, thereby reducing the impact of randomness. The Monte Carlo method is unbiased as it receives the actual rewards. However it suffers from high variance since it relies on complete episodes, i.e., the fully constructed formulas.
In [19], the actor-critic architecture is adopted, with an actor network as the policy and a critic network for reducing the variance in the training process. Naturally, the seminal Proximal Policy Optimization (PPO) [21] was chosen to train the networks. This paper proposes that the actor-critic architecture and PPO may not be suitable for this factor mining framework. Specifically, the aforementioned alpha factors mining framework has trajectory feedback, i.e., non-zero reward can only be obtained after completing a full trajectory [22, 23]. This makes the critic network of PPO difficult to extract useful training signals from the intermediate states, leading to inevitable biases when attempting to estimate the value of the state and thus slow converging speed. Furthermore, the critic network typically has the same scale with the policy network, meaning that two large networks need to be updated in each iteration of PPO. This not only makes the training harder, but leads to doubled parameter updating time[24].
Based on the above discussions, this paper proposes a new DRL-based method for formulaic alpha factors mining, named QuantFactor REINFORCE (QFR). QFR abandons the critic network of AlphaGen and train the policy using the REINFORCE[25], a seminal Monte Carlo method for policy gradient estimation. In this regard, the biases of PPO can be vanished. The high variance of REINFORCE is mitigated in two aspects. For the variance caused by the environment, we observe that the underlying MDPs has deterministic transitions, which satisfies the Dirac distribution. That is, once a new action is selected, the new state is uniquely determined by the current sequence and the newly generated token within the RPN representation. This ensures that REINFORCE will not be degenerated by the environmental variance. For the variance caused by the policy gradient estimation, QFR employs a greedy policy to generate a novel baseline, which has been theoretically shown to bound the variance and leads to a lower variance than the original REINFORCE. Additionally, we introduce the Information Ratio (IR) for reward shaping to better balance returns and risks. As a result, QFR enjoys faster convergence speed to the optimal policy and produces steadier alpha factor signals across various market volatility.
@@
To evaluate our proposed QFR algorithm, we conducted extensive experiments on multiple real-world asset datasets. Our experimental results show that the set of formulaic alpha factors generated by the QFR algorithm outperforms those generated by previous methods. It is also verified that theoretical results properly predict trends in the experimental results, even encountering distinct market volatility.
@@
  • We propose a more stable and efficient reinforcement learning algorithm for mining formulaic alpha factors. Unlike prior research that employs the actor-critic framework, we argue to discard the critic network due to the trajectory feedback and theoretically show that the high variance due to the absence of the critic network can be significantly mitigated.
  • Two new components are proposed for DRL-based formulaic alpha factors mining. A novel baseline that employs a greedy policy is proposed to reduce the variance of the policy gradient estimation. Additionally, IR is introduced as a reward shaping mechanism to balance returns and risks. We also provide a set of theoretical results, including an analysis of the training variance under state transition functions with various distributions, the derivation of the upper bound of the variance after introducing the baseline, and the proof that the variance decreases compared to REINFORCE.
@@ II-A Automatic Mining for Alpha Factors
The work in [19] is the first to employ DRL for formulaic alpha factors mining. It employs MDPs to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using reinforcement learning. Compared to works [34, 35, 36] that directly use reinforcement learning to build trading agents, this framework uses historical quantitative and price data of assets as input to find a set of formulaic alpha factors with strong interpretability, avoiding the black-box problem and overcoming the limitations of [11, 12, 13, 14, 15] in independently mining individual formulaic factors, such as homogenization among factors, lack of synergy, and difficulty in adapting to dynamic market changes. The characteristics of reinforcement learning, supervised learning, tree models, and heuristic algorithms when applied to mining alpha factors are shown in Table I.
@@
Reinforcement Learning Supervised Learning Tree Models Genetic Programming
@@
II-B The REINFORCE algorithms
III Problem Formulation and Preliminaries
This section first introduces the definition of formulaic alpha factors and their sequences with RPN. Next, the MDPs for mining the formulaic alpha factors modeled by Yu et al. [19] are detailed. Lastly, the two seminal methods for solving the modeled MDPs, i.e., PPO and REINFORCE, are compared to motivate the proposed QFR algorithm.
III-A Alpha Factors for Predicting Asset Prices
Consider a real market with $n$ assets over $L$ trading days. On each trading day $l\in\{1,2,\cdots,L\}$, each asset $i$ corresponds to a feature vector $\mathbf{X}_{li}\in\mathbb{R}^{m\times d}$. This vector consists of $m$ raw market features, such as open, high, low, close, and volume values, over the recent $d$ days. Here, $\mathbf{x}_{lij}\in\mathbb{R}^{d\times 1}$ denotes the sequence of values for the $j$-th raw feature over these $d$ days.
Next, we define an alpha factor function $f$, which transforms the feature matrix for all $n$ assets on the $l$-th trading day, represented as $\mathbf{X}_{l}=[\mathbf{X}_{l1},\mathbf{X}_{l2},\cdots,\mathbf{X}_{ln}]^{\mathsf{T}}\in\mathbb{R}^{n\times m\times d}$, into alpha factor values $\mathbf{z}_{l}=f(\mathbf{X}_{l})\in\mathbb{R}^{n\times 1}$. Specifically, $\mathbf{z}_{l}$ holds the alpha factor values for all assets on the $l$-th trading day. The real asset feature dataset over $L$ days is denoted as $\mathcal{X}=\{\mathbf{X}_{l}\}$.
After the policy model outputs a new formulaic alpha factor, this alpha will be added to an alpha factors pool $\mathcal{F}$, say =$\mathcal{F}=\{f_{1},f_{2},...,f_{K}\}$ with $K$ factors. The linear combination model [51] is adopted to compute the asset price predicted by the alpha factor pool[19]. Specifically, suppose each $k$-th alpha factor ($1\leq k\leq K$) is associated with a weight $w_{k}$, and the asset price is calculated as $\mathbf{z}_{l}^{\prime}=\sum^{K}_{k=1}w_{k}f_{k}(\mathbf{X}_{l})$. In other words, the mining of alpha factors is to search a set of alpha factors and use them in combination. The weight of each factor indicates its exposure on the assets and is optimized using gradient descent. The loss function of learning the weights vector $\mathbf{\omega}\in\mathbb{R}^{K\times 1}$ is defined as the mean squared error (MSE) between the model output and the ground-truth asset prices $\mathcal{Y}=\{\mathbf{y}_{l}\}$ with $l\in\{1,2,\cdots,L\}$ and $\mathbf{y}_{l}\in\mathbb{R}^{n\times 1}$:
$L(\mathbf{\omega})=\frac{1}{L}\sum_{l=1}^{L}\left\|\mathbf{z}_{l}^{\prime}-\mathbf{y}_{l}\right\|^{2},$ (1)
Due to significant differences in the scales of different alpha factor values, they are first normalized to have a mean of 00 and a maximum value of $1$. If the number of factors in the factor pool exceeds a certain threshold, the alpha factor with the smallest weight and its corresponding weight will be discarded.
Figure 1: Three uniquely interchangeable forms of an alpha factor: (a) formulaic expression; (b) tree structure; (c) RPN sequence.
@@ III-A Alpha Factors for Predicting Asset Prices
III-B Formulaic Alpha Factors
Formulaic alpha factors are mathematical expressions that can be represented using RPN, which is a sequence of tokens. Tokens include various operators, the original volume-price features, fundamental features, time deltas, constants, and sequence indicators. The operators include elementary functions that operate on single-day data, known as cross-sectional operators (e.g., Abs($x$) for the absolute value $|x|$, Log($x$) for the natural logarithm log($x$)), as well as functions that operate on a series of daily data, known as time-series operators (e.g., Ref($x,l$) for the expression $x$ evaluated at $l$ days before the current day, where $l$ denotes a time token, such as 10d (10 days)). The Begin (BEG) token and Separator (SEP) token of the RPN representation are used to mark the beginning and end of the sequence. Table III illustrates a selection of these tokens as examples. Such formulas can naturally be represented by an expression tree, with each non-leaf node representing an operator, and the children of a node representing the original volume-price features, fundamental features, time deltas, and constants being operated on. Each such expression has a unique post-order traversal, using RPN. An example of a formulaic alpha expression, together with its corresponding tree and RPN sequence, is shown in Fig. 1. More examples of formulaic alpha expressions derived from the well-known Alpha101 [37] and their corresponding RPN sequences are shown in Table II.
TABLE III: Examples of Formulaic Tokens
Token Types Token Instances
Operators Abs($x$), Log($x$), Ref($x,l$)
Features open, high, low, close
Time Deltas 10d, 20d, 50d
Constants -10, -5, -0.01, 0.01, 5, 10
Sequence Indicators BEG, SEP
To measure the effectiveness of a formulaic alpha factor, the Pearson correlation coefficient between the ground-truth asset price $\mathbf{y}_{l}$ and the combination factor value $\mathbf{z}_{l}^{\prime}$, also known as the Information Coefficient (IC), is commonly employed as the performance indicators which calculates as follows:
$IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)=\frac{\text{Cov}(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l})}{\sigma_{\mathbf{z}^{\prime}_{l}}\sigma_{\mathbf{y}_{l}}}.$ (2)
where $\mathrm{Cov}(\cdot,\cdot)$ indicates the covariance matrix between two vectors, and $\sigma_{\cdot}$ means the standard deviation of a vector. The averaged IC values over all $L$ trading days are denoted as $\overline{IC}=\mathbb{E}_{l}\left[IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)\right]=\frac{1}{L}\sum^{L}_{l=1}IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)$.
III-C MDPs for Mining Formulaic Alpha Factors
@@
The process of generating a sequence of tokens that can equivalently represent a formulaic alpha factor with RPN is modeled as an MDP[19], which can be described in the classic tuple of $\{\mathcal{S},\mathcal{A},P,r\}$. Specifically, $\mathcal{A}$ denotes the finite action space, consisting of a finite set of candidate tokens as actions $a$. $\mathcal{S}$ represents the finite state space, where each state at the $t$-th time step corresponds to the sequence of selected tokens, representing the currently generated part of the formulaic expression in RPN, denoted as $\mathbf{s}_{t}=\mathbf{a}_{1:t-1}=[a_{1},a_{2},\cdots,a_{t-1}]^{\mathsf{T}}$. Our goal is to train a parameterized policy $\pi_{\mathbf{\theta}}:\mathcal{S}\rightarrow\mathcal{A}$, which generates optimal formulaic alpha factors by iteratively selecting from the candidate tokens. The selecting process can thus be modeled as $a_{t}\sim\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{a}_{1:t-1}\right)$, where the action $a_{t}$ is the next token following the currently generated part of the expression $\mathbf{a}_{1:t-1}$ in RPN sequence.
The transition function $P$ defines the state transitions. When $\mathbf{a}_{1:t-1}$ and $a_{t}$ are already known, then $\mathbf{s}_{t+1}=\mathbf{a}_{1:t}$ is uniquely determined, which means that the state transition function $P$ satisfies the Dirac distribution:
$P\left(\mathbf{s}_{t+1}\mid\mathbf{a}_{1:t}\right)=\left\{\begin{array}[]{ll}1&\text{ if }\mathbf{s}_{t+1}=\mathbf{a}_{1:t}\\ 0&\text{ otherwise. }\end{array}\right.$ (3)
An legal formulaic always starts with the begin (token) BEG, followed by any token selected from $\mathcal{A}$, and ends when the separator token (SEP) is selected or the maximum length is reached. Obviously, any generated sequence cannot be guaranteed to be a legal RPN sequence, therefore [19] only allow specific actions to be selected in certain states to ensure the correct format of the RPN sequence. For more details about these settings, please refer to [19].
@@
The reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns values to the state-action pairs and is set to $r\left(\mathbf{a}_{1:T}\right)=\overline{IC}$ in [19]. The optimization objective in this MDP is to learn a policy $\pi_{\mathbf{\theta}}$ that maximizes the expected cumulative reward over time:
$J(\theta)=\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\mathbf{\theta}}}\left[r\left(\mathbf{a}_{1:T}\right)\right].$ (4)
@@
It is clear that non-zero rewards are only received at the final $T$-th step, which evaluates the quality of a complete formulaic factor expression, not individual tokens:
$r\left(\mathbf{s}_{t},a_{t}\right)=\left\{\begin{array}[]{ll}0&\text{ if }t\neq T\\ r\left(\mathbf{a}_{1:T}\right)&\text{ otherwise. }\end{array}\right.$ (5)
@@
Expressions that are syntactically correct might still fail to evaluate due to the restrictions imposed by certain operators. For example, the logarithm operator token is not applicable to negative values. Such invalidity can not be directly detected. Therefore, these expressions are assigned a reward of $-1$ (the minimum value of IC) to discourage the policy from generating these expressions.
III-D Comparing REINFORCE with PPO
Gradient ascent is a typical way of learning the policy $\pi_{\mathbf{\theta}}$ by iteratively optimizing $\theta$: $\mathbf{\theta}_{k+1}\leftarrow\mathbf{\theta}_{k}+\eta_{k}\cdot g\left(\mathbf{\theta}_{k}\right)$, where $g(\mathbf{\theta}_{k})$ is the policy gradient at the $k$-th iteration and $\eta_{k}$ represents the corresponding learning rate. The general policy gradient $g(\mathbf{\theta})$ is calculated as follows: